# The SOFIA Massive (SOMA) Radio Survey. II. Radio Emission from High-Luminosity Protostars

### v1

@@ 5 Summary and Conclusions
Many of our fainter radio sources have radio luminosities that are consistent with a simple power law extrapolation from a lower-mass YSO sample (Anglada, 1995). However, this radio flux is about ten times higher than a theoretical model of radio emission from shock ionization of a fiducial TCA massive protostar (Gardiner et al., 2024). Nevertheless, shock ionization appears to be the most likely explanation for many of these sources, especially given the morphology of extended jet knots, while a contribution from photoionization, especially at the Inner scale, cannot be excluded. However, in the high radio luminosity regime, photoionization is expected to be dominant. Overall, these data are important for constraining theoretical models of massive protostar ionization and how this links to protostellar structure and evolutionary state.
6 Acknowledgments
F. S. M. acknowledges support from the NRAO NINE (National and International Non-traditional Exchange) program. V. R. acknowledges support from NSF grant AST–2206437. J.C.T. acknowledges support from NSF grant AST–2206437 and ERC Advanced Grant 788829 (MSTAR). R.F. acknowledges support from the grants PID2023-146295NB-I00, and from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033 and by “European Union NextGenerationEU/PRTR”. The National Radio Astronomy Observatory and Green Bank Observatory are facilities of the U.S. National Science Foundation operated under cooperative agreement by Associated Universities, Inc. We acknowledge Simon Purser and Andres Guzmán for providing the ATCA images for sources G339.88–1.26 and IRAS 16562–3959, respectively.
VLA, ATCA
@@ Appendix B Evolutionary tracks from radio luminosities
Figure B1 makes a comparison of the SOMA Radio data at the inner (top) and SOMA (bottom) scale, with the full set of protostellar evolutionary tracks from Zhang & Tan (2018), i.e., with 10, 30, 60, 120, 240 and 480 with 0.316 g cm-2 (left), 1 g cm-2 (center) and 3.16 g cm-2 (right). In these models the radio luminosities are not set from radiation transfer like the tracks from (Tanaka et al., 2016), but instead assumes a simple spherical HII region. Meanwhile, Figure B2 shows the same information but with the average of the good models of the isotropic bolometric luminosity (Table C.1 of SOMA V).

### v3

@@ 5 Summary and Conclusions
Many of our fainter radio sources have radio luminosities that are consistent with a simple power law extrapolation from a lower-mass YSO sample (G. Anglada, 1995). However, this radio flux is about ten times higher than a theoretical model of radio emission from shock ionization of a fiducial TCA massive protostar (E. C. Gardiner et al., 2024). Nevertheless, shock ionization appears to be the most likely explanation for many of these sources, especially given the morphology of extended jet knots, while a contribution from photoionization, especially at the Inner scale, cannot be excluded. However, in the high radio luminosity regime, photoionization is expected to be dominant. Overall, these data are important for constraining theoretical models of massive protostar ionization and how this links to protostellar structure and evolutionary state.
@@ Appendix B Evolutionary tracks from radio luminosities
Figure B1 makes a comparison of the SOMA Radio data at the inner (top) and SOMA (bottom) scale, with the full set of protostellar evolutionary tracks from Y. Zhang & J. C. Tan (2018), i.e., with 10, 30, 60, 120, 240 and 480 with 0.316 g cm-2 (left), 1 g cm-2 (center) and 3.16 g cm-2 (right). In these models the radio luminosities are not set from radiation transfer like the tracks from (K. E. I. Tanaka et al., 2016), but instead assumes a simple spherical HII region. Meanwhile, Figure B2 shows the same information but with the average of the good models of the isotropic bolometric luminosity (Table C.1 of SOMA V).